OpenStudy (anonymous):

suppose that a 3 digit number randomly generated using digits from 1-8. Find the probability that the number has at least one repeated digit

OpenStudy (anonymous):

if two household are independently selected, what is the probability that both have income in the $30,000- $39,999 range? Income 0- $9,999 6% $10,000-$19,999 9% $20,000- 29,999 13% $30,000-$39,999 28% $40,000-$49,999 18% $50,000-$59,999 17% $60,000 or more 9%

OpenStudy (dan815):

1-prob of no repeated digits

OpenStudy (dan815):

1-(8*7*6/8^3)

OpenStudy (anonymous):

You can count the ways to have at least two repeated digits: 8 ways to have exactly 3 repeated digits (111, 222 etc.) plus 8*7 ways to have only the first two digits repeated (112, 113 etc.) plus 8*7 ways to have only the first and the last digits repeated (121, 131 etc.) plus 8*7 ways to have only the last two digits repeated (211, 311 etc.).

OpenStudy (anonymous):

wait i wrote it wrong the number one it will be in total 6 digits (1,2,3,5,7,8) so i did this 6.5.4 / 6.6.6 =120/ 216 i simplify = 5/6 = 0.833

OpenStudy (anonymous):

wait if forgot to -1 = 0.167?